Find the relation of the variable w if the growth rate is given by


dw/dt=(kw(α-w))/α, k>0, at t=0,w=α/(1+β)


1
Expert's answer
2021-08-23T07:28:20-0400

Let us find the relation of the variable "w" if the growth rate is given by


"\frac{dw}{dt}=\frac{kw(α-w)}{α},\ \ k>0," at "t=0,\ w=\frac{α}{1+β}."


It follows that


"\frac{\alpha dw}{w(α-w)}=kdt"


"\int\frac{\alpha dw}{w(α-w)}=\int kdt"


"\int(\frac{1}{α-w}+\frac{1}{w})dw=\int kdt"


"\ln|w|-\ln|\alpha -w|=kt+C_1"


"\ln|\frac{w}{\alpha -w}|=kt+C_1"


"\frac{w}{\alpha -w}=C_2e^{kt}"


"w=C_2e^{kt}(\alpha -w)"


"w=C_2e^{kt}\alpha -wC_2e^{kt}"


"w (1+C_2e^{kt})=C_2e^{kt}\alpha"


"w =\frac{C_2\alpha e^{kt}}{1+C_2e^{kt}}"


Since at "t=0,\ w=\frac{α}{1+β}," we have that


"\frac{α}{1+β}=\frac{C_2\alpha}{1+C_2}"


"\frac{1}{1+β}=\frac{C_2}{1+C_2}"


"1+C_2=C_2+\beta C_2"


"1=\beta C_2"


"C_2=\frac{1}{\beta}"


We conclude that


"w =\frac{\frac{1}{\beta}\alpha e^{kt}}{1+\frac{1}{\beta}e^{kt}}=\frac{\alpha e^{kt}}{\beta+e^{kt}}."